Please read these instructions before posting any event on Fermilab Indico

Indico will be unavailable due to server patching on Wed , June 19th, 7 a.m. to 7:30 a.m. Central Time

ND-GAr: HPgTPC+ECAL Weekly Meeting

US/Central
Disappearance Room (WH13SE) and Zoom

Disappearance Room (WH13SE) and Zoom

    • 10:00 AM 10:20 AM
      Roundtable Updates 20m
      Speakers: Alfons Weber (JG-Uni Mainz & Fermilab), Diego Gonzalez Diaz, Naseem Khan, Tanaz Mohayai (Fermilab)

      GOAT:

      (Tanaz)  Still trying to figure out what happened to the IROC.  Can't give voltages above 2150 V anymore.  Moving on to GEM tests.  Don't need to make any changes to flanges.    Will test with 55Fe at 1 atm soon, and later move on to high-pressure tests.  The detectors are are 3 GEM stacks (with its own divider)  with an additional cathode on top.  Diego notes that the resistor config might need to be optimized for pressure and gas type.    Tend to put lower gain on the last one.  Diego also wonders if there are enough power supplies to power each electrode separately without a resistor divider.    These GEMs have been run before by the company in Ar/CO2.  The transfer fields depend on E/P but the multiplication inside the doesn't scale the same way (usually scales less).    The relative voltages likely don't scale together.  Is it safe to use the existing divider chain?  Diego suggests maybe setting up a meeting with the CERN group.  GOAT does have enough HV channels, and has 4 HV feedthroughs (need drift, and then one for each GEM layer) so maybe enough.  Can maybe add one or two more if needed.   Pablo can maybe run simulations as a starting point. 

      (Alan)  Something happened ALICE production IROC when running at higher voltages at 7 atm.  Eventually realized that pads were floating.  They are now grounded, but can't get above 2150 V (regardless of pressure).  Don't see much obviously wrong with it or evidence of arcing.  But there is discoloration on the pad plane, and looks worse than another production chamber that was in storage.  Discoloration on vias and near where the source was. Wire planes look very good and electronics are still ok.  Diego says that you lose can insulation in the pad plane, so suspect that the breakdown might be in the planes.   Could also maybe be a feedthrough issue.

       

      TOAD:

      (Naseem)  Been doing vacuum tests and rates of rise tests, using both the backing pump and also the turbo.  Seems to be working.   Currently vessel is at vacuum, with pumps off, and will leave it like that during collab meeting.  Also resolved an issue with the slow control pressure not agreeing with the display on the Pirani pressure sensor.  For DAQ, fake data is now going through the system.  Some changes made to DUNE DAQ and hope to get TOAD format fully integrated soon.  Will hopefully put gas lines in this week and work is underway to be able to run the hydraulic pump.  But things are moving along.  Maybe ready for next round of ORCs after collab meeting.

      (Ioannis) In London waiting for frontend cards.  There are 5 that were at Pittsburgh, shipped to FNAL.    If things look good then produce all 150 cards.

      (Anezka) Working on the event display where there is an offset in the drift direction, but event display should be readyish after that.

      Alan asks if OROC has been run with ungrounded pads in the past.  It did come with the copper shims, so the other pads were likely grounded. 

       

      Light collection:

      (Diego)  Will give a report next week.

       

      ECAL:

      (Vivek)  Still working to make some productions to test code that Leo pushed recently.  Hopes to start maybe int he next week. 

      (Leo) Working on it again.  Wants to tweak up the clustering parameter and check the track to ecal cluster matching.  Also wants to look again for electron neutrino interactions.   New geometry doesn't have as much material upstream (not cylindrically symmetric anymore), and this made it hard to see nue at lower momentum due to conversions in upstream material.   Vivek thinks that the geometry is still 12-sided and symmetric, but Leo will check.